OpenStudy (anonymous):

lim {tan(pi/4+x)} whole raised to 1/x x->0 . please find the limit ? waiting eagerly !

OpenStudy (anonymous):

do u know the LHR?

OpenStudy (anonymous):

Hopital i mean

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok\[y=\left( \tan(x+\frac{\pi}{4})\right)^{\frac{1}{x}}\]\[\ln y= \frac{\ln(\tan(x+\frac{\pi}{4}))}{x}\]use LHR to find the limit of RHS of later equation then find y

OpenStudy (anonymous):

:D